OpenStudy (anonymous):

What is the vertex of the graph of y = 3x2 + 2x + 1? A.-1/3, 0 B.-1/3, 2/3 C.1/3, 2 D.2/3, 11/3

OpenStudy (anonymous):

first coordinate of the vertex of \(y=ax^2+bx+c\) \[-\frac{b}{2a}\] which in your case is \[-\frac{2}{3\times 2}=-\frac{1}{3}\]